OpenStudy (anonymous):

Expand the logarithm Quesition about placement - My question is do I need to add parenthesis to logarithm answers?

OpenStudy (anonymous):

My book has these all over the place so I'm not sure.

OpenStudy (anonymous):

I think you should add parenthesis idk i might be wrong

OpenStudy (jdoe0001):

the "2" is multiplying the logarithmic function, thus if the logarithm expands to the "sum" version, the "2" has to multiply the expanded logarithmic version then

OpenStudy (anonymous):

so then yes add the parenthesis?

OpenStudy (jdoe0001):

\(\bf log_4(3xyz)^2\implies 2\left( log_43+log_4x+log_4y+log_4z \right) \\ \quad \\ 2log_43+2log_4x+2log_4y+2log_4z\)

OpenStudy (anonymous):

okay you're confusing me you just wrote both down

OpenStudy (jdoe0001):

hmmm nope, the last line is the parenthesized version, expanded

OpenStudy (anonymous):

okay so I don't add parenthesis.

OpenStudy (jdoe0001):

well.. distributing the "2" will be, expanding, thus

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

I'm confused now and idk what I'm doing

jimthompson5910 (jim_thompson5910):

the first step is to pull down the exponent 2 using this rule \[\Large \log_{b}(x^y) = y*\log_b(x)\]

OpenStudy (anonymous):

right

jimthompson5910 (jim_thompson5910):

after you pulled down the 2, you will have \[\Large 2\log_{4}\left(3xyz\right)\]

jimthompson5910 (jim_thompson5910):

then you'll use the rule log(x*y) = log(x) + log(y) to expand out log(3xyz)

OpenStudy (anonymous):

yep

jimthompson5910 (jim_thompson5910):

use the rule log(x*y) = log(x) + log(y) to get... \[\Large 2\log_{4}\left(3xyz\right)\] \[\Large 2[\log_{4}\left(3xyz\right)]\] \[\Large 2[\log_{4}(3)+\log_{4}(xyz)]\] \[\Large 2[\log_{4}(3)+\log_{4}(x)+\log_{4}(yz)]\] \[\Large 2[\log_{4}(3)+\log_{4}(x)+\log_{4}(y)+\log_{4}(z)]\]

jimthompson5910 (jim_thompson5910):

I used the rule in the smallest pieces possible. So it's a bit more dragged out than it has to be. You can just go from log(3xyz) to log(3)+log(x)+log(y)+log(z) in one step really

jimthompson5910 (jim_thompson5910):

anyways, the point is that 2 on the outside is being multiplied by each term inside. You can leave the 2 out there like it is, but make sure you have surrounding parenthesis around those log terms don't write \[\Large 2\log_{4}(3)+\log_{4}(x)+\log_{4}(y)+\log_{4}(z)\] instead write \[\Large 2[\log_{4}(3)+\log_{4}(x)+\log_{4}(y)+\log_{4}(z)]\] or write \[\Large 2(\log_{4}(3)+\log_{4}(x)+\log_{4}(y)+\log_{4}(z))\]

OpenStudy (anonymous):

yeah I have options to choose from and the parenthesis around the 3,x,y,and z isn't an option it's just 2log4 3+log4 x+log4 y +log4 z or 2(log4 3+log4 x+log4 y +log4 z) so I feel like I will just go w/ the parenthesis.

jimthompson5910 (jim_thompson5910):

then go with 2(log4 3+log4 x+log4 y +log4 z)

OpenStudy (anonymous):

my book has examples written like yours so i think that's why I was confused. Thank you for helping me

jimthompson5910 (jim_thompson5910):

as for what jdoe0001 did, he simply used the distribution rule to multiply the '2' by each term inside eg: 2*(x+y) = 2*x + 2*y

jimthompson5910 (jim_thompson5910):

I'm glad I could help clear things up